# CHaRNet: Conditioned Heatmap Regression for Robust Dental Landmark Localization

### v4

@@ Abstract
Identifying anatomical landmarks in 3D dental models is vital for orthodontic treatment, yet manual placement is complex and time-consuming. Although some machine learning approaches have been proposed for automatic tooth landmark detection in 3D Intraoral Scans (IOS), none provide a fully end-to-end solution that bypasses teeth segmentation, limiting practical applicability.
We introduce CHaRNet (Conditioned Heatmap Regression Network), the first fully end-to-end deep learning framework for tooth landmark detection in 3D IOS. Unlike traditional two-stage workflows that segment teeth before detecting landmarks, CHaRNet directly operates on the input point cloud, thus reducing complexity and computational overhead. Our method integrates four modules: (1) a point cloud encoder, (2) a point cloud decoder with a heatmap regression head, (3) a teeth presence classification head, and (4) the novel Conditioned Heatmap Regression (CHaR) module. By leveraging teeth presence classification, the CHaR module dynamically adapts to missing teeth and enhances detection accuracy in complex dental models.
We evaluate CHaRNet using five point cloud learning algorithms on a clinical dataset of 1,214 annotated 3D models. Both the dataset and code will be publicly released [^1] Resources will be made public upon acceptance to address the lack of open datasets in orthodontics and inspire further research. CHaRNet achieves a Mean Euclidean Distance Error (MEDE) of 0.51 mm on typical dental models and 1.28 mm across all dentition types, with corresponding Mean Success Rates (MSR) of 87.06% and 82.40%, respectively. Notably, it exhibits robust performance on irregular geometries, including models with missing teeth. This end-to-end approach streamlines orthodontic workflows, enhances 3D IOS analysis precision, and supports efficient computer-assisted treatment planning.
@@ 1 Introduction
Automatic 3D anatomical keypoints or landmarks detection plays a crucial role in computer-aided orthodontics, allowing tasks such as orthodontic planning, prosthetic design, and the diagnosis of dental anomalies [1]. Intraoral scans (IOS) are commonly utilized to capture accurate digital surface models of dentition. These scanners represent the 3D surface of teeth, typically as point clouds or meshes. These 3D dental models are highly valuable for simulating procedures such as tooth extraction, movement, deletion, and rearrangement, allowing dentists to predict treatment outcomes more effectively. As a result, digital teeth models have the potential to streamline dentists’ workflows and reduce the time spent on labor-intensive tasks.
Existing approaches to perform 3D dental tasks such as teeth segmentation and landmark detection often rely on point cloud-based learning methods [2, 3, 4, 5, 6]. Although effective for general point cloud problems, these methods struggle in dental applications, where incomplete geometries and subtle variations in anatomy, such as missing teeth or the presence of third molars, lack specialized handling [7, 3, 8, 9]. Furthermore, the absolute lack of publicly available dental datasets impedes progress by preventing fair comparisons between methods and limiting accessibility to this field to researchers outside the specialized domains [8, 9].
@@
    We introduce CHaRNet, an end-to-end deep learning (DL) model that directly detects landmarks on the IOS point cloud without relying on previous teeth segmentation, drastically reducing the latency of the overall computation.
@@
    We propose the Conditioned Heatmap Regression (CHaR) module, a novel enhancement to point cloud learning architectures, which dynamically adapts to missing teeth.
@@
    We construct and make publicly available a dataset of $1,214$ annotated 3D digital teeth models, which includes diverse complexities, detailed reference landmarks and a taxonomy of dentition types to evaluate models across varying levels of complexity.
Our results highlight the effectiveness of the proposed CHaR module, demonstrating significant performance improvements across multiple point cloud learning models. Notably, the CHaR-augmented PointMLP, referred to as CHaRNet, achieves a Mean Euclidean Distance Error (MEDE) of 0.51 mm on typical dental models, defined as those with a complete set of teeth excluding third molars, and 1.28 mm across all dentition types. It achieves a Mean Success Rate (MSR) of 87.06% on typical dental models and 82.40% across all dentition types. These findings underscore the robustness of CHaRNet, particularly in challenging cases involving several missing teeth.
The remainder of this paper is organized as follows: Related works are presented in Section 2. Section 3 describes the dataset construction process, including the proposed taxonomy and data preprocessing steps. Section 4 formally introduces the problem and the proposed method, including the CHaR module and its integration into point cloud learning architectures. Section 5 presents a thorough evaluation of the proposed approach across different point cloud architectures. Section 6 discusses the broader implications, limitations, and potential directions for future work. Finally, Section 7 summarizes the key contributions and results of this work.
@@ 2.1 Point cloud learning
Traditionally, three main stream methods have been used for 3D point cloud processing: projection-based, voxel-based and point cloud-based. Projection-based methods try to firstly project the original point cloud to a simpler 2D domain to later apply common 2D operations for feature extraction [10, 11, 12, 13]. While simple and effective, the projection operations over point sets inherently collapse potentially useful geometric information. Voxel-based methods transform the irregular nature of point clouds to regular 3D voxels (the equivalent of pixels in 3D) then apply 3D convolution operations [14, 15]. Given the sparse nature of 3D data, directly applying 3D convolutions over the voxel input results in a highly inefficient computation. To take advantage of this sparsity, sparse convolutions [16, 17, 18] greatly reduce computation and memory requirements by evaluating the convolution operation only in occupied voxels. Finally, point cloud-based methods directly process point clouds as unordered point sets. PointNet [19] pioneered this approach by directly processing point sets using shared multilayer perceptrons (MLP). PointNet++ [20] improves its predecessor by hierarchically applying set abstraction operations. PointNeXt [21] further improves PointNet++ by using modern training strategies such as data augmentation, new optimization techniques, and increased model size. The authors of PointMLP [22] hypothesize that local geometrical information may not be the key to point cloud learning and introduce a simple and pure residual MLP network that performs competitively with more complicated methods. DGCNN [23] and similar graph-based approaches [24, 25, 26] perform typical graph operations like message passing over the point cloud previously connected as some form of graph. Given the success of the self-attention mechanism, PointTransformer [27] designed self-attention layers for point cloud processing and use them to perform tasks such as scene semantic segmentation or object classification in 3D scenes. Since then several studies have kept using transformers [28, 29, 30, 31, 32].
@@ 2.2 Landmark detection
In recent years, landmark localization task has become popular in medical image analysis [55, 56, 57, 58]. However, only a handful of works have been proposed for landmark detection in 3D dental models. [1] proposed a set of analytical approaches to identify dental-specific features (e.g., cusps, marginal ridges, and grooves) on digital dental meshes. At individual tooth level, [3] designed a novel neural network architecture for the joint tasks of predicting landmarks and axes. [59] presented a method for landmark detection in 3D dental meshes that leverages a multi-view approach, transforming the task into the 2D domain. In this way, the network detects landmarks through heatmap regression across multiple viewpoints. ALR [60] first determines the orientation of the scan, then uses local maxima along the vertical axis as an initial approximation for landmarks. This is followed by analyzing surface gradient and curvature information to identify the shape and boundaries of each tooth. c-SCN [7] was proposed as an end-to-end method for tooth segmentation and landmark localization but only on teeth crowns. TS-MDL [8] proposed a two-stage framework to subsequently perform teeth segmentation and landmark detection on IOS. By harnessing the segmentation outputs, authors used a variant of PointNet to detect landmarks for each individual tooth. While effective, this approach comes at a high computational cost derived from separating the process into two stages. Besides, the used dataset includes mainly regular dental models without considering complex dental models with missing teeth and one or two third molars. ALIIOS algorithm [9] combines image processing, segmentation, and DL to identify dental landmarks on IOS by synthesizing 3D patches with the output of a 2D U-Net. However, most of these methods rely on a previous segmentation of teeth, which adds considerable latency to the overall system.
@@
Ours (2025) Yes Yes Upper and lower Yes $1,214$
@@
3 Data construction
Existing datasets for landmark detection in 3D intraoral scans are very small, private and do not include frequent anomalies of the dentition. A comparison between these datasets is provided in Table 1. In this work, we construct a comprehensive dataset of 1,214 digital 3D dental models derived from post-processed raw scans. Each model is annotated with detailed reference landmarks and categorized into a taxonomy of dentition types to facilitate the evaluation of automatic 3D landmark detection systems under varying levels of difficulty. Our dataset does not only provide a robust foundation for training and evaluating point cloud learning methods but also addresses a significant gap in the orthodontic field, where publicly available annotated datasets are scarce. In this regard, we make our dataset publicly available so that the community can explore new research ideas on it. This is also an attempt to encourage the orthodontic community to publish their datasets and hence allow for a fair comparison between different AI methods, attracting more practitioners to the field of computer-assisted orthodontics, and encouraging the development of new domain-specific technologies.
To enable accurate and consistent annotations, each dentition in our dataset contains five landmarks per tooth, making up to 80 landmarks in total (depending on the number of teeth present in each dentition). These landmarks represent key points for subsequent tasks, such as treatment planning, and are illustrated in Figure 1 and detailed in Table 2. Specifically, for each tooth, the five landmarks identify important anatomical structures that help guide orthodontic measurements and interventions. This meticulous labeling ensures that every model in the dataset captures essential dental characteristics, ranging from normal anatomical variations to more complex anomalies, thereby broadening the scope and applicability of the dataset for diverse research and clinical needs.
@@ 3 Data construction
Figure 2: Examples of each dentition type based on our proposed taxonomy. The rows represent the presence or absence of third molars: the first row (0) indicates no third molars are present, while the second row (1) indicates the presence of at least one third molar. The columns correspond to the number of missing teeth in the dentition.
@@ 3.1 Dentition layout and taxonomy
Table 4: Dentition taxonomy and dataset count.
@@
00 No third molar, no missing teeth 668
@@ 3.3 Data preprocessing
In order to feed the data to the point cloud learning methods, our 3D dental models, composed of $67,000$ points and $130,000$ faces on average, are downsampled to $10,000$ vertices, as it allows enough resolution for the point cloud models to learn the desired mapping while reducing memory and computation costs. Some original and downsampled point clouds are shown in Figure 4.
@@ 3.4 Ground-truth generation
[…]
@@ 4 Methodology
In this section, we first formally define our problem in Section 4.1. Then, our proposed framework is introduced in Section 4.2. Finally, our implementation details are described in Section 4.3.
@@ 4.1 Problem statement
To train the model, we employ two loss functions: one for the landmark localization task and another for the presence/absence classification task.
@@
4.2 Network
Figure 6: The pipeline of our proposed method: Given an input point cloud of a 3D dental model, (1) the point cloud encoder extracts the internal features, then these features are fed into the next two stages simultaneously. (2) The initial landmark regression stage and (3) the presence classification of teeth are performed in two separate branches. Finally, (4) the conditioned heatmap regression module makes use of the presence classification to adjust the initial landmark regression.
In this paper, we propose CHaRNet (Conditioned Heatmap Regression Network), an end-to-end DL-based landmark detection method for 3D dental models, which does not rely on previous segmentation of teeth. It is able to handle missing teeth and fasten inference, making it suitable for clinical applications. Our proposed method is shown in Figure 6, where (1) the point cloud encoder extracts the internal representation, (2) the point cloud decoder plus the heatmap regression head performs the initial heatmap-based regression of landmarks, (3) the classification head predicts the presence or abscense of every teeth, and (4) our novel conditioned heatmap regression module adjusts every heatmap’s scores based on the presence classification of its associated teeth.
@@ 4.2 Network
The point cloud encoder is in charge of extracting a reduced internal representation of the whole input point cloud. Point cloud learning methods usually perform this dimensionality reduction and feature selection hierarchically through set abstraction operations [PointNet2, PointNext]. This learned high-level features will subsequently be fed to the decoder plus heatmap regression head, and the teeth presence classification.
@@
In general, when a tooth is missing, the vanilla landmark regression network is not explicitly designed to map all landmarks associated with that tooth to the null point. To address this issue, the conditioned heatmap regression (CHaR) module adjusts the initial landmark regression to take into account the predicted presence probability of each tooth. For clarity, each landmark $k$ is denoted using a tuple of two digits $(t,g)$, where the first digit $t=1,2,...,16$ indicates the tooth associated with the landmark, and the second digit $g=1,2,...,5$ indicates the specific landmark within the group of 5 landmarks associated with tooth $t$. For example, landmark $16$ (the 1st landmark of the 4th tooth) is denoted as $(4,1)$ in this notation.
@@
4.3 Implementation details
Dataset splitting
We follow the typical three-way holdout setup for model comparison and performance estimation [61], that is, the 1,214 dental models are divided into train (70%), validation (15%) and test (15%) sets. The training set is used to adjust the weights of the neural network, the validation set is used to adjust the network and training hyper-parameters, and the test set is used to estimate model performance on previously unseen data. Given that the 1,214 dental models contain maxilar and mandibular samples from 923 patients, we split the patients into three subsets and subsequently assign their corresponding dental models to avoid any correlation between them.
Training details
All base models and CHaR versions in the results section were trained by minimizing the MSE loss (Formula 3) and the combined loss (Formula 5), respectively, using the Adam optimizer [62] for a total of 100 epochs using a batch size of 16. The initial learning rate is set to 0.005 and is progressively reduced using the cosine learning rate decay scheduler. To prevent overfitting, weight decay was set to 0.003, serving as a regularization term. Additionally, dropout rates were tuned between 0.3 and 0.5 across different model configurations to enhance generalization. Pytorch [63] was used as the DL framework to conduct all experiments. With respect to hyper-parameters controlling the relative importance in the final loss value in the CHaR versions, $\lambda_{\text{reg}}$ and $\lambda_{\text{cls}}$ are set to 0.001 and 1 respectively, given the natural bigger quantity of the landmark localization loss.

### v5

# CHaRM: Conditioned Heatmap Regression Methodology for Accurate and Fast Dental Landmark Localization

@@ Abstract
Identifying anatomical landmarks in 3D dental models is essential for orthodontic treatment, yet manual placement is labor-intensive and requires expert knowledge. While machine learning methods have been proposed for automatic landmark detection in 3D Intraoral Scans (IOS), none provide a fully end-to-end solution that avoids costly tooth segmentation.
We present CHaRM (Conditioned Heatmap Regression Methodology), the first fully end-to-end deep learning approach for tooth landmark detection in 3D IOS. CHaRM integrates four components: a point cloud encoder, a decoder with a heatmap regression head, a teeth-presence classification head, and the novel CHaR module. The CHaR module leverages teeth-presence information to adapt to missing teeth, improving detection accuracy in complex dental cases. Unlike two-stage workflows that segment teeth before landmarking, CHaRM operates directly on IOS point clouds, reducing complexity, avoiding error propagation, and lowering computational cost.
We evaluated CHaRM with five point cloud learning backbones on IOSLandmarks-1k, a new dataset of 1,214 annotated 3D dental models. Both the dataset and code will be publicly released [^1] Resources will be made available upon acceptance to address the scarcity of open data in orthodontics and foster reproducible research.
CHaRM with PointMLP, named CHaRNet, achieved the best accuracy and efficiency. Compared to state-of-the-art methods (TSMDL and ALIIOS), CHaRNet reduced mean Euclidean distance error to 0.56 mm on standard dental models and 1.12 mm across all dentition type, while delivering up to 14.8× faster inference on GPU. This end-to-end approach streamlines orthodontic workflows, enhances the precision of 3D IOS analysis, and enables efficient computer-assisted treatment planning.
@@ 1 Introduction
Automatic 3D anatomical keypoints or landmarks detection plays a crucial role in computer-aided orthodontics, allowing tasks such as orthodontic planning, prosthetic design, and the diagnosis of dental anomalies [1]. Intraoral scans (IOS) are commonly utilized to capture accurate digital surface models of dentitions. These scanners represent the 3D surface of teeth, typically as point clouds or meshes. These digital representations are invaluable for simulating procedures such as tooth extraction, movement, deletion, and rearrangement, allowing dentists to predict treatment outcomes more effectively. Consequently, digital dental models have the potential to streamline dentists’ workflows and reduce the time spent on labor-intensive tasks.
Existing approaches to perform 3D dental tasks such as teeth segmentation and landmark detection often rely on point cloud-based learning methods [2, 3, 4, 5, 6]. Although effective for general point cloud problems, these methods often struggle in dental applications due to the specialized nature of dental anatomies, which frequently exhibit incomplete geometries and subtle variations, such as missing teeth or the presence of third molars [7, 3, 8, 9]. Moreover, traditional landmark detection workflows typically involve a preliminary tooth segmentation stage [8, 9]. While widely adopted in digital dentistry, teeth segmentation introduces significant drawbacks. It is computationally demanding, requires careful post-processing to ensure stable contours, and most critically, can propagate segmentation errors into subsequent landmark detection methods, potentially compromising accuracy. In many clinical scenarios, such as prosthetic design or simple orthodontic assessments, explicit tooth segmentation may not be clinically necessary, yet current pipelines mandate its inclusion, adding unnecessary overhead.
Another significant limitation in this field is the scarcity of publicly available dental datasets. This restricts progress by preventing fair comparisons among methods and limiting accessibility for researchers from outside specialized dental domains [8, 9].
@@
    We propose the Conditioned Heatmap Regression (CHaR) module, specifically designed for point cloud learning architectures. This module dynamically adapts landmark detection to account for missing teeth, significantly improving accuracy across complex dental anatomies.
@@
    We introduce CHaRM, an end-to-end Deep Learning (DL) methodology that directly detects landmarks on the IOS point cloud without relying on previous teeth segmentation, drastically reducing the latency of the overall computation. Concretely, CHaRNet, CHaRM based on PointMLP as point cloud processing model, is more accurate and several orders of magnitude faster than the state-of-the art (SOTA) TSMDL [8] and ALIIOS [9], in both CPU and GPU.
@@
    We construct and make publicly available IOSLandmarks-1k, a dataset of $1,214$ 3D digital teeth models with tooth segmentation and landmarks annotations, with the hope to foster broader research and benchmarking. IOSLandmarks-1k includes diverse complexities and a taxonomy of dentition types to evaluate models across varying levels of complexity.
Our results highlight the effectiveness of the proposed CHaRM and CHaR module, demonstrating significant performance improvements across multiple point cloud learning models. Notably, CHarNet achieves a Mean Euclidean Distance Error (MEDE) of 0.56 mm on standard dental models, defined as those with a complete set of teeth excluding third molars, and 1.12 mm macro-averaging across all dentition types and a Mean Success Ratio (MSR) of 85.2% and 68.5%, respectively, surpassing previous SOTA methods. Furthermore, by eliminating the need for previous teeth segmentation, CHaRNet speeds up inference time several orders of magnitude. These findings underscore the robustness and efficiency of CHaRNet, particularly in challenging cases involving missing teeth and/or third molars.
The remainder of this paper is organized as follows: Section 2 presents related works. Section 3 describes the dataset construction process, including the proposed taxonomy and data preprocessing steps. Section 4 formally introduces the problem and the proposed CHaRM, detailing the CHaR module and its integration with point cloud learning architectures. Section 5 presents a thorough evaluation of the proposed approach across different point cloud architectures and the comparison with SOTA methods. Section 6 discusses the broader implications, limitations, and potential directions for future work. Finally, Section 7 summarizes the key contributions and results of this work.
@@ 2 Related works
This section reviews related work in two areas: point cloud learning methods (Section 2.1) and landmark detection approaches (Section 2.2).
@@ 2.1 Point cloud learning
At present, three main stream methods have been used for 3D point cloud processing: projection-based, voxel-based and point cloud-based. Projection-based methods try to firstly project the original point cloud to a simpler 2D domain to later apply common 2D operations for feature extraction [10, 11, 12, 13]. While simple and effective, the projection operations over point sets inherently collapse potentially useful geometric information. Voxel-based methods transform the irregular nature of point clouds to regular 3D voxels (the equivalent of pixels in 3D) then apply 3D convolution operations [14, 15]. Given the sparse nature of 3D data, directly applying 3D convolutions over the voxel input results in a highly inefficient computation. To take advantage of this sparsity, sparse convolutions [16, 17, 18] greatly reduce computation and memory requirements by evaluating the convolution operation only in occupied voxels. Finally, point cloud-based methods directly process point clouds as unordered point sets. PointNet [19] pioneered this approach by directly processing point sets using shared multilayer perceptrons (MLP). PointNet++ [20] improves its predecessor by hierarchically applying set abstraction operations. PointNeXt [21] further improves PointNet++ by using modern training strategies such as data augmentation, new optimization techniques, and increased model size. The authors of PointMLP [22] hypothesize that local geometrical information may not be the key to point cloud learning and introduce a simple and pure residual MLP network that performs competitively with more complicated methods. DGCNN [23] and similar graph-based approaches [24, 25, 26] perform typical graph operations like message passing over the point cloud previously connected as some form of graph. Given the success of the self-attention mechanism, PointTransformer [27] designed self-attention layers for point cloud processing and use them to perform tasks such as scene semantic segmentation or object classification in 3D scenes. Since then several studies have kept using transformers [28, 29, 30, 31, 32].
@@ 2.2 Landmark detection
In recent years, landmark localization task has become popular in medical image analysis [55, 56, 57, 58]. However, only a handful of works have been proposed for landmark detection in 3D dental models. [1] proposed a set of analytical approaches to identify dental-specific features (e.g., cusps, marginal ridges, and grooves) on digital dental meshes. At the individual tooth level, [3] designed a novel neural network architecture for the joint tasks of predicting landmarks and axes. [59] presented a method for landmark detection in 3D dental meshes that leverages a multi-view approach, transforming the task into the 2D domain. In this way, the network detects landmarks through heatmap regression across multiple viewpoints. ALR [60] first determines the orientation of the scan, then uses local maxima along the vertical axis as an initial approximation for landmarks. This is followed by analyzing surface gradient and curvature information to identify the shape and boundaries of each tooth. c-SCN [7] was proposed as an end-to-end method for tooth segmentation and landmark localization but only on teeth crowns. TS-MDL [8] proposed a two-stage framework to subsequently perform teeth segmentation and landmark detection on IOS. By harnessing the segmentation outputs, authors used a variant of PointNet to detect landmarks for each individual tooth. While effective, this approach comes at a high computational cost derived from separating the process into two stages. Besides, the dataset used includes mainly regular dental models without considering complex dental models with missing teeth and one or two third molars. ALIIOS algorithm [9] combines image processing, segmentation, and DL to identify dental landmarks on IOS by synthesizing 3D patches with the output of a 2D U-Net. However, most of these methods rely on a previous segmentation of teeth, which adds considerable latency to the overall system.
@@
Teeth3DS+ [61] (2024) Yes Yes Upper and lower Yes $340$
IOSLandmarks-1k (2025) Yes Yes Upper and lower Yes $1,214$
@@
3 IOSLandmarks-1k
Existing datasets for landmark detection in 3D intraoral scans are very small, private and/or do not include frequent anomalies of the dentition. A comparison between these datasets is provided in Table 1. Teeth3DS+ [61] is the only publicly available dataset. However, it only contains 340 annotated samples, which may be insufficient for training modern neural networks. In this work, we construct IOSLandmarks-1k, a comprehensive dataset of 1,214 digital 3D dental models derived from post-processed raw scans. The post-processing step involves removing common artifacts found in raw scans, resulting in standardized dental models that retain only clinically relevant information. Each model is annotated with detailed reference landmarks and categorized into a taxonomy of dentition types to facilitate the evaluation of automatic 3D landmark detection systems under varying levels of difficulty. Furthermore, teeth segmentation annotations are also provided in order to also train and evaluate methods that require previous segmentation like TSMDL and ALIIOS (see Figure 2). IOSLandmarks-1k does not only provide a robust foundation for training and evaluating teeth segmentation and landmark detection methods, but also addresses a significant gap in the orthodontic field, where publicly available annotated datasets are scarce. In this regard, we make IOSLandmarks-1k publicly available so that the community can explore new research ideas on it. This is also an attempt to encourage the orthodontic community to publish their datasets and hence allow for a fair comparison between different AI methods, attracting more practitioners to the field of computer-assisted orthodontics and encouraging the development of new domain-specific technologies. Full details on the segmentation and landmark annotations format will be provided in OSLandmarks-1k public repository.
To enable accurate and consistent annotations, each dentition in our dataset contains five landmarks per tooth, making up to 80 landmarks in total (depending on the number of teeth present in each dentition). A visual illustration is provided in Figure 1 and details in Table 2. These landmarks represent key points for subsequent tasks, such as treatment planning. Specifically, for each tooth, the five landmarks identify important anatomical structures that help guide orthodontic measurements and interventions. This meticulous labeling ensures that every model in the dataset captures essential dental characteristics, ranging from normal anatomical variations to more complex anomalies, thereby broadening the scope and applicability of the dataset for diverse research and clinical needs.
@@ 3 IOSLandmarks-1k
Figure 2: Examples of each dentition type based on our proposed taxonomy. The rows represent the presence or absence of third molars: the first row (0) indicates no third molars are present, while the second row (1) indicates the presence of at least one third molar. The columns correspond to the number of missing teeth in the dentition. Each tooth is colored with its associated color given a predefined color map.
@@ 3.1 Dentition layout and taxonomy
Table 4: Description and number of samples of IOSLandmarks-1k’s dentition types.
@@
$0$ No third molar, no missing teeth 668
@@ 3.3 Data preprocessing
In order to feed the data to the point cloud learning methods, our 3D dental models, composed of $67,000$ points and $130,000$ faces on average, are downsampled to $10,000$ vertices, as it allows enough resolution for the point cloud models to learn the desired mapping while reducing memory and computation requirements. Some original and downsampled point clouds are shown in Figure 4.
@@ 3.4 Ground-truth generation
where $d$ is the Euclidean distance function, and $\sigma$ is a hyperparameter representing the standard deviation of the Gaussian distance field, which we set to $2\text{mm}$.
@@ 4 Methodology
In this section, we first formally define our problem in Section 4.1. Then, our proposed framework is introduced in Section 4.2.
@@ 4.1 Problem statement
To train the model, we employ two loss functions: one for the landmark localization task and another for the tooth presence/absence classification task.
@@
4.2 CHaRM
Figure 6: Pipeline of CHaRM. Given an input point cloud of a 3D dental model, (1) the point cloud encoder extracts internal features, which are then simultaneously processed by two parallel branches: (2) a landmark regression branch that generates initial heatmaps for each landmark, and (3) a teeth presence classification branch that predicts the existence of each tooth. Finally, (4) the Conditioned Heatmap Regression (CHaR) module uses the presence predictions to refine the initial landmark heatmaps. Point cloud encoder and decoder (shown in red) can be any point cloud learning architecture, such as PointMLP or PointTransformer.
In this paper, we propose CHaRM (Conditioned Heatmap Regression Methodology), an end-to-end DL-based landmark detection methodology for 3D dental models, which does not rely on previous segmentation of teeth. It is able to handle missing teeth and fasten inference, making it suitable for clinical applications. Our proposed methodology is shown in Figure 6, where (1) the point cloud encoder extracts the internal representation, (2) the point cloud decoder plus the heatmap regression head performs the initial heatmap-based regression of landmarks, (3) the classification head predicts the presence or absence of every tooth, and (4) our novel conditioned heatmap regression module adjusts every heatmap’s scores based on the presence classification of its associated teeth.
@@ 4.2 CHaRM
The point cloud encoder is in charge of extracting a reduced internal representation of the whole input point cloud. Point cloud learning methods usually perform this dimensionality reduction and feature selection hierarchically through set abstraction operations [20, 21]. These learned high-level features will subsequently be fed to the decoder plus heatmap regression head and the teeth presence classification.
@@
In general, when a tooth is missing, the vanilla landmark regression network is not explicitly designed to map all landmarks associated with that tooth to the null point. To address this issue, the conditioned heatmap regression (CHaR) module adjusts the initial landmark regression to take into account the predicted presence probability of each tooth. For clarity, each landmark $k$ is denoted using a tuple of two digits $(t,g)$, where the first digit t=1,2,,16t=1,2,...,16 indicates the tooth associated with the landmark, and the second digit g=1,2,,5g=1,2,...,5 indicates the specific landmark within the group of 5 landmarks associated with tooth $t$. For example, landmark $16$ (the 1st landmark of the 4th tooth) is denoted as $(4,1)$ in this notation.